OpenStudy (anonymous):

what condition must a and b satisfy for matrix [a+b b-a a-b b+a] to be orthogonal? I'm desperate to find the solution, i don't understand all of this

OpenStudy (badhi):

$$Q=\left[\begin{matrix} a+b&b-a\\ a-b&b+a\end{matrix}\right]$$ $$Q^T=\left[\begin{matrix} a+b&a-b\\ b-a&b+a\end{matrix}\right]$$ to become orthogonal, $$QQ^T=Q^TQ=I$$( http://en.wikipedia.org/wiki/Orthogonal_matrix) $$\begin{align*}QQ^T&=\left[\begin{matrix} a+b&b-a\\ a-b&b+a\end{matrix}\right]\left[\begin{matrix} a+b&a-b\\ b-a&b+a\end{matrix}\right]\\&=\left[\begin{matrix} (a+b)^2+(b-a)^2 & (a^2-b^2)+(b^2-a^2)\\ (a^2-b^2)+(b^2-a^2)&(a-b)^2+(b+a)^2\end{matrix}\right]\\ &=\left((a+b)^2+(b-a)^2 \right)\left[\begin{matrix} 1&0\\ 0&1\end{matrix}\right]\end{align*}$$ for this to become I,$$(a+b)^2+(b-a)^2=1\implies 2a^2+2b^2=1$$

OpenStudy (anonymous):

thank you! i understand